# Enabling Fine-Grained Operating Points for Black-Box LLMs

### v1

@@ A.4 Additional Experiments with Simple Prompting Techniques to Extract Verbalized Probabilities
Tables 3 and 4 compare the different prompting techniques along with the proposed Ours-Unsup method that uses additive noise on both the 11 individual datasets and the joint dataset. We report the averaged metrics for the 11 datasets results. While the tailored prompts do increase the cardinality compared to Prompt-Naive, all of them still have very low cardinalities compared to Ours-Unsup. The increase in cardinality also does not necessarily reflect in improved granularities, limiting their utility. Predicting multiple outputs and choosing one of them randomly (Multiple-Pred) results in the best performance in terms of operational granularity but results in degredation of PRAUC and ROC-AUC, particularly on the individual datasets. Note that due to the extremely low cardinality of these methods, the PRAUC and ROC-AUC tend to be over-estimates due to linear interpolation between points.
@@ A.9 Extended Experiment Results with The Proposed Approaches
In this section, we provide additional measurements that correspond to the experiment results presented in the main body. Figure 11 extends Figure 4 that reported PRAUCs with $2\sigma$ error bars for the individual dataset (11ds) and joint dataset (joint) settings by providing AUROCs. Figures 12 and 13 show PR and ROC curves across 5 seeds, extending the 5Figure 5 (a): from the main body that focused on precision-recall curves of the benchmark and the proposed methods using Claude on a single seed. The optimization in the learning based approach Ours-Sup-1call can sometimes fail due to an incorrect choice of learning rate combined with extremely small dataset sizes (refer Figure 12(i)). We ignore such runs (and corresponding runs from other methods) when calculating the aggregated metrics. Finally, the Figures 14, 15, 16 and 17 breaks down the analysis provided for dataset-specific training and testing covered in Figure 4 (Top) into performance metrics on individual datasets, providing PR and ROC AUCs.
@@
Figure 12 (a): Claude Seed #1
Figure 12 (b): Nova Seed #1
Figure 12 (c): Claude Seed #2
Figure 12 (d): Nova Seed #2
Figure 12 (e): Claude Seed #3
Figure 12 (f): Nova Seed #3
Figure 12 (g): Claude Seed #4
Figure 12 (h): Nova Seed #4
Figure 12 (i): Claude Seed #5
Figure 12 (j): Nova Seed #5
@@
Figure 13 (a): Claude Seed #1
Figure 13 (b): Nova Seed #1
Figure 13 (c): Claude Seed #2
Figure 13 (d): Nova Seed #2
Figure 13 (e): Claude Seed #3
Figure 13 (f): Nova Seed #3
Figure 13 (g): Claude Seed #4
Figure 13 (h): Nova Seed #4
Figure 13 (i): Claude Seed #5
Figure 13 (j): Nova Seed #5

### v2

@@ A.4 Additional Experiments with Simple Prompting Techniques to Extract Verbalized Probabilities
Tables 5 and 6 compare the different prompting techniques along with the proposed Ours-Unsup method that uses additive noise on both the 11 individual datasets and the joint dataset. We report the averaged metrics for the 11 datasets results. While the tailored prompts do increase the cardinality compared to Prompt-Naive, all of them still have very low cardinalities compared to Ours-Unsup. The increase in cardinality also does not necessarily reflect in improved granularities, limiting their utility. Predicting multiple outputs and choosing one of them randomly (Multiple-Pred) results in the best performance in terms of operational granularity but results in degredation of PRAUC and ROC-AUC, particularly on the individual datasets. Note that due to the extremely low cardinality of these methods, the PRAUC and ROC-AUC tend to be over-estimates due to linear interpolation between points.
@@ A.9 Extended Experiment Results with The Proposed Approaches
In this section, we provide additional measurements that correspond to the experiment results presented in the main body. Figure 11 extends Figure 4 that reported PRAUCs with $2\sigma$ error bars for the individual dataset (11ds) and joint dataset (joint) settings by providing AUROCs. Figures 12 and 13 show PR and ROC curves across 5 seeds, extending the 5 from the main body that focused on precision-recall curves of the benchmark and the proposed methods using Claude on a single seed. The optimization in the learning based approach Ours-Sup-1call can sometimes fail due to an incorrect choice of learning rate combined with extremely small dataset sizes (refer Figure 12(i)). We ignore such runs (and corresponding runs from other methods) when calculating the aggregated metrics. Finally, the Figures 14, 15, 16 and 17 breaks down the analysis provided for dataset-specific training and testing covered in Figure 4 (Top) into performance metrics on individual datasets, providing PR and ROC AUCs.
@@
(a) Claude Seed #1
(b) Nova Seed #1
(c) Claude Seed #2
(d) Nova Seed #2
(e) Claude Seed #3
(f) Nova Seed #3
(g) Claude Seed #4
(h) Nova Seed #4
(i) Claude Seed #5
(j) Nova Seed #5
@@
(a) Claude Seed #1
(b) Nova Seed #1
(c) Claude Seed #2
(d) Nova Seed #2
(e) Claude Seed #3
(f) Nova Seed #3
(g) Claude Seed #4
(h) Nova Seed #4
(i) Claude Seed #5
(j) Nova Seed #5